OpenStudy (anonymous):

Find the value of x: -x/3 = 4/9 x = -133 x = 133 x = 1.33 x = -1.33

OpenStudy (anonymous):

-x/3 = 4/9 multiply each side by 3 -x=1.3333... divide each by -1 x=-1.333333....